OpenStudy (kmays7):

The legs of a right triangle measure 4 and 6 inches. The area of the triangle is

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

If you let the side with length 4 be the base, then the side with length 6 is the height. If you let the side with length 6 be the base, then the side with length 4 is the height. That means that either way, you have the base and the height.

OpenStudy (2500161):

A = ½ * b * h We let: b = 4 h = 6 So.. A = ½ * 4 * 6 = 12 inches²

OpenStudy (mathstudent55):

Do you know the formula of the area of a triangle?

OpenStudy (mathstudent55):

\(\Large A_{triangle} = \dfrac{bh}{2} \)